OpenStudy (anonymous):

Solve the quadratic equation x^2+3x-4=0 and round to the nearest hundredth when necessary

OpenStudy (anonymous):

Well, you can use the quadratic formula.

OpenStudy (anonymous):

Or factoring.

OpenStudy (mrnood):

factorising will giove you the answer very quickly you need to get 2 numbers a & b a+b = 3 a * b = -4 then (x-a)(x-b) = x^2 + 3x -4 so (x-a) =0 OR (x-b) = 0 those are the 2 values for x that solve the equation

OpenStudy (anonymous):

\[x^2+3x-4=0\\x^2+(4-1)x-4=0\\x^2+4x-1x-4=0\\x(x+4)-1(x+4)=0\\(x+4)[x-1]=0\\ \implies\ (x+4)=0\ OR\ (x-1)=0\\ \implies\ x=-4\ OR\ x=1\]